Prev Next

SW&IT Meeting Minutes 2022-05-16

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220502
LR20220502 - Bec, Matthieu 80.0
LR20220502 - Mechtley, Brandon 72.0
LR20220502 - Eychaner, Glenn 68.0
LR20220502 - Edmeades, David 76.0
LR20220502 - Grenz, Paul 12.0
LR20220502 - Hooper, Stephen 0.0
LR20220502 - Zhang, Yang 77.0
LR20220502 - Zhang, Xianyu 89.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management, AOSWSYS, Luci-gui-WG
    • MG support contract amends
    • shark-vis talk
    • JS software and computer discussions
    • safety training
  • This week
    • ARGOS github synch (Lorenzo?)
    • AzCam kickoff with M. Lesser
    • SOUL comm meeting & report

YZ

  • Previous:
    • TMS software discussion and re-analysis of key algorithms (on-going work).
    • MODS/LBC issues discussion with Olga.
    • Some new MODS/LBC software work proposed by Olga (will discuss them).
    • Some investigation into adding API support in lbto.dms Python module (maybe not the right place for archive new images?).
    • LBC code study.
    • Completed safety training.
  • Next:
    • Continue some TMS discussions and general work (software/wiki).
    • Continue various MODS/LBC issues/work discussion with Olga (including planning for the new proposed software).
    • Continue LBC code study.

DE

  • Last week
    • IssueTrak bug
      • They found the bug--related to a 20-character limit on entity names in text email. No reason for that, so they removed it.
      • Pushed fix to our site already
    • LBTI ingestion
      • On track, mostly holding for their run
      • Have a couple of test tasks this week
    • Backup server sundries
      • Installed the (very basic) PERC monitoring software
      • Ginned up a script to parse the machine-unfriendly output and give regular status updates
      • Couple of refinements to go. When it's sorted I'll change the target email to IT and suppress the "all good" messages
    • Accounts add on ssh.mountain
      • 1/3 response rate so far
  • This week
    • Ketiv babysitting for some maintenance task; not sure why it's not just done in the background. I will find out.
    • Said LBTI testing/prep
    • Script tweaks
    • Get on the SSD schedule to lock down LeRoy 's time
    • Quick look at new IT API, but the support person I talked to re: last ticket didn't think that it was yet scoped for ticket work.
    • Doc appointment Friday afternoon

PSG

  • Last week
    • Phasecam debugging, testing.
    • LMIRCAM debugging testig.
    • LBTI software shakedown and testing in general.
    • SHARK-NIR software fixes and updates.
  • This week
    • AzCam kickoff with M. Lesser
    • SHARK-NIR software additions and fixes.

XZ

  • Last week
    • study and create the MMC actuator list;
    • analyze the MMC test daytime test;
    • support LBTI DX daytime work and two nights;
    • switch DX ASM configuration for day and night;
    • analyze unexpected failure on SX ASM
    • NCPA for Shark-NIR
    • test and analysis strange behavior of ring acuators on DX ASM
  • This week
    • test and analysis strange behavior of ring acuators on DX ASM
    • analyze unexpected failure on SX ASM
    • NCPA for Shark-NIR

GE

  • Last week:
    • Released LUCI 2.14.12:
      • Catches the Ice Timeout exceptions found in IT 8573.
      • Includes final fixes for IT 8619 (problem centering slit).
    • Added code to LUCI to track Journalizer connections for IT 8529 (Observer GUI stops updating). Not yet deployed.
    • Tested IT 8326 (Pixel Histogram). Could not replicate problem, but added my own notes on Aladin UI oddities.
      • During that testing, found that Aladin still contains some internal deadlocks, which could cause 8355. Investigating.
      • In the process, refreshed my memory on using the Java Console and Debugger to monitor running Java processes. (The process has to be started with monitoring and/or debugging enabled in the first place, which adds a small amount of extra overhead; the live GUIs are not started with monitoring and debugging enabled, but we might want to change that.)
    • Second meeting of the LUCI GUI Working Group.
  • This week:

BM

  • Last week
    • set_z0 - no problems
    • Closed IT 8602, moving on from 8556 (interrupted offsets) but will let others close (maybe more to discuss re: skip frames)
    • SetNCPA discussion/save fits file
    • Investigating new offset (IT 8635), ocam boot problems.
    • VM testing (setncpa, RTDB emtpy string issue, gain ramp)
  • This week
    • Some test support for Juan Carlos (changing rotation/moving baysides), for IT 8635
    • Put together test version on sxadsec for Shark SetNCPA if they decide to use it ..
    • Need to upgrade ao-supervisor "stable" on all the machines. Will include everything from last weekend's DEC tests except the BCU power up fix (IT 8553)
    • Finish this gain ramp testing finally ...
    • ocam boot issues re: 8553
    • soul commissioning SW report--anything specific we should be looking for?

SH

  • Meetings
  • UITS, Tucson network
  • Room 507 computers
  • 64-bit Perl Tk
  • OVMS telemetry
  • SSL certs
  • AWS, UITS, Windows Servers

next:

  • Vacation
  • Core switches
  • solidworker
Topic revision: r10 - 16 May 2022, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback